8. Ok, this one drives me crazy. I know I have to set all equations to zero... this word problem needs to be solved and written out with factoring. What does that look like here? What is the solution? I'm very confused!
Thank you for your help!!


The area of a rectangle is 24 square meters. Find the length and width of the rectangle if its length is 2 meters greater than its width. Use an equation and the formula ® area of a rectangle = (width)(length).

Answer by nerdybill(7384)

The area of a rectangle is 24 square meters. Find the length and width of the rectangle if its length is 2 meters greater than its width.
.
Let w = width of rectangle
then from: "its length is 2 meters greater than its width" we get
w+2 = length of rectangle
.
We know that the area of any rectangle is: (length * width) AND
the problem tells us that the area is 24 sq meters.
From which we derive our equation:
w(w+2) = 24
.
w^2 + 2w = 24
w^2 + 2w - 24 = 0
factoring we get:
(w+6)(w-4) = 0
Solutions for w={-6,4}
We can throw out the negative answer because it doesn't make sense.
.
Therefore:
width = 4 meters
length = w+2 = 4+2 = 6 meters